A market consists of two individuals . Their demand equations are:

P= 4 - 0.25Qand P = 10 - 0.5Q2

a) What is the market demand equation?

b) At a price of $2, what is the point price elasticity for each person and for the market?



1
Expert's answer
2022-04-04T09:15:08-0400

"P_1=4-0.25Q_1"

"P_2=10-0.5Q_2"

A) The market demand, Q, is given as the total of all individual demands:

"Q_1=16-4P_1"

At P=$2, Q1=8

"Q_2=20-2P_2"

At P=$2, Q2= 16

"\therefore Q=36-6P"


B) "\epsilon=\frac{{d}{Q}}{{d}{P}}×\frac{P}{Q}"

Point elasticity of the market at Q=24 and P=2

"\epsilon_m=-6×\frac{2}{24}"

"\epsilon_m=-0.17"


"\epsilon_{P{_{1}}}=-4×\frac{2}{8}"

"\epsilon_{P_{1}}=-1"


"\epsilon_{P_{2}}=-2×\frac{2}{16}"

"\epsilon_{P_{2}}=-0.25"